At a given instant, a 3.5. A current flows in the
Answer #1

I= C dV/dt

Capacitance of the capacitor is

C = εoA/d = 8.85x10-12 (0.0142) /(0.014) =1.239x10-14F (Square plates area is a2 = 0.0142 and d = 0.014)

dV/dt = I/C = 3.5/1.239x10-14 =2.83x1014V/s

the electric field in a parallel plate capacitor is

dE/dt = (dV/dt )/ d = 2.83x1014 / 0.014 = 2.02x1016 V/ ms